OpenStudy (anonymous):

The value -1 is a lower bound for the zeros of the function shown below. f(x)=x^4+x^3-11x^2-9x+18 A. True B. False

OpenStudy (anonymous):

can we use calculus?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

FALSE x=-3,-2,1,3 are roots of polynomial i was trying to find a proof. i found all of roots!!!!!!:D by letting x=-(t+1) u will have \[t^4+3t^2-8t^2-12t+16=t^4-8t^2+16+3t^3-12t=(t^2-4)^2+3t(t^2-4)\\=(t^2-4)(t^2+3t-4)=(t-2)(t+2)(t-4)(t+1)\] Another way u can prove it by Mean Value Theorem f(-1)>0 and f(-2.5)<0 so there is a root on the interval (-2.5 , -1)

OpenStudy (anonymous):

oh ok thank you!! :)

OpenStudy (anonymous):

welcome dear:D